A simple random sample of 12 e-readers of a certain type had the following minutes of battery life.

287, 311, 262, 392, 313, 291, 350, 316, 286, 339, 314, 291

Assume that it is reasonable to believe that the population is approximately normal and the population standard deviation is 54. What is the upper bound of the 95% confidence interval for the battery life for all e-readers of this type?

Round your answer to one decimal places (for example: 319.4). Write only a number as your answer. Do not write any units.
Answer #1

Solution

The sample mean is

Mean = (x / n) )


=287+311+262+392+313+291+350+316+286+339+314+291/12

=3752/12

=312.6667

Mean = 312.7

= 8.8

n = 12

At 95% confidence level the z is ,

  = 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.960

Margin of error = E = Z/2* (/n)

= 1.960 * (54/ 12 )

=30.5

At 95% confidence interval estimate of the population mean is,

+ E

312.7 + 30.5

= 343.2

The upper bound = 343.2